OpenStudy (kittiwitti1):

Prove the identity arcsin(x)+arcos(x)=π/2

OpenStudy (kittiwitti1):

LaTeX:\[\text{Prove the identity }\sin^{-1}{x}+\cos^{-1}{x}=\frac{\pi}{2}\]

zepdrix (zepdrix):

Ooo this problem looks really familiar :d Hmm thinking...

zepdrix (zepdrix):

Remember that your trig functions take an angle as in input, \(\large\rm \sin(\theta)\) and output something else (usually a ratio). Our inverse function does the opposite, right? It's taking in this argument x, and it's outputting an angle. So let's assign a variable to this arcsin(x), and we'll use a variable that used often for an angle.\[\large\rm \arcsin(x)=\theta\]

OpenStudy (kittiwitti1):

\(x\)=coordinate of a trig fcn?

zepdrix (zepdrix):

No, let's try to avoid thinking of x in that way. :) Mainly just because it's the argument of BOTH arcsine and arccosine.

zepdrix (zepdrix):

Using this variable assignment, we can write it in terms of sine,\[\large\rm \large\rm \arcsin(x)=\theta\qquad\to\qquad \sin(\theta)=x\]

zepdrix (zepdrix):

The next step is a little tricky. You have to recall that sine and cosine are `co-functions`. That means that they have this relationship,\[\large\rm \sin(\theta)=\cos(90-\theta)\]

OpenStudy (kittiwitti1):

._. ... Enjoy this while I unscramble my brain.

zepdrix (zepdrix):

Haha that's awesome XD

zepdrix (zepdrix):

These steps getting sort of confusing? Hmm I'm trying to think if there is another approach... I'm not sure.

OpenStudy (kittiwitti1):

Okay, my thought process is still disjointed. I can't connect the dots with your explanation... though I may be tired, and that's why

zepdrix (zepdrix):

I'll leave the rest of the steps if you want to at least see them. Maybe something will click later on.

OpenStudy (kittiwitti1):

😨 okay;

OpenStudy (kittiwitti1):

the original format of the question is with -1 exponents, not arc(fcn)

zepdrix (zepdrix):

So we have our variable assignment for theta, and we rewrote the relationship in terms of the sine function,\[\large\rm \sin(\theta)=x\]We'll write this in terms of the cosine function using what we know about co-functions,\[\large\rm \cos(90-\theta)=x\]We don't really want degrees for this problem though, so I'll write it in a more appropriate way,\[\large\rm \cos\left(\frac{\pi}{2}-\theta\right)=x\]From here we'll switch to inverse cosine, swapping the arguments,\[\large\rm \arccos(x)=\frac{\pi}{2}-\theta\]

zepdrix (zepdrix):

Just something to keep in mind with trig functions, -1 exponent NEVER refers to a power, similar to function notation.\[\large\rm f^{-1}(x)\ne\frac{1}{f(x)}\]\[\large\rm \sin^{-1}(x)\ne\frac{1}{\sin(x)}\]That -1 is a special power, it's always reserved for the inverse function. Sometimes we write -1 power, sometimes arc, sall good :)

OpenStudy (kittiwitti1):

yes I understand this 👌👌

zepdrix (zepdrix):

We started with this relationship for theta,\[\large\rm \arcsin(x)=\theta\]And now we've found another relationship for theta,\[\large\rm \arccos(x)=\frac{\pi}{2}-\theta\]

OpenStudy (kittiwitti1):

\[\arcsin{x}=\sin^{-1}{x}\]yes?

zepdrix (zepdrix):

Yes, good.

OpenStudy (kittiwitti1):

ok making sure... think I got it so far

zepdrix (zepdrix):

So we started with this,\[\large\rm \arcsin(x)+\arccos(x)\]replacing each of these with something involving this angle theta,\[\large\rm = \theta+\left(\frac{\pi}{2}-\theta\right)\]

OpenStudy (kittiwitti1):

AH, I finally got what you were saying. bleh am slow today

zepdrix (zepdrix):

I'm curious to see what the other guy has to say XD Maybe he has a much simpler method haha

OpenStudy (caozeyuan):

 If we assume x=b/c then arcsin of x is just theta and arccos of x is just the other acute angle,and they must add up to 90 or pi/2 becuase they are in right angle trianlge

OpenStudy (kittiwitti1):

*brain is scrambled again*

OpenStudy (kittiwitti1):

uh... I think I need to stick to one method of explaining. brain does not compute 2 levels of logic 😂😂

zepdrix (zepdrix):

Ya that triangle gives some clarity to this whole co-function idea.